OpenStudy (anonymous):

Can someone help me find the Rate Of Change in 5x+3y =-2?

OpenStudy (amistre64):

rate of change tends to be a calculus term; are you dong calculus or just trying to find the slope of a line?

OpenStudy (anonymous):

Find the slope of the line.

OpenStudy (amistre64):

site seems to be a little on the fritz today :/

OpenStudy (amistre64):

one way to find the slope is to let x =0 to find y then let x=1 and find y again; then subtract the x=1 value of y from the x=0 value of y

OpenStudy (anonymous):

3y= - 5x -2 Y= - 5/3x - 2/3 If x=0, y= - 2/3 (OY intercept) m= - 5/3 (the slope)

OpenStudy (amistre64):

slope can be defined as the distance and direction the value of y moves as x moves by 1

OpenStudy (anonymous):

5x + 3y = -2 y = mx + b Subtract 5x from each side 5x + 3y - 5x = 3y -2 - 5x = -2 -5x 3y = -2 -5x Divide each item by 3 3y/3 = -2/3 -5x/3 y = -5x/3 -2/3

OpenStudy (anonymous):

3y= - 5x -2 Y= - 5/3x - 2/3 If x=0, y= - 2/3 (OY intercept) m= - 5/3 (the slope)